# Bayesian evidence for two peaks in the sound speed in cold dense QCD

### v1

@@ I Introduction
It can be understood in terms of the geometric interpretation of eq. 8, namely $\Delta P$ as the area in fig. 1. Continuations of NS EOSs satisfying eq. 10 must have segments above the mean value EOS. Had the sound speed of such high-density models were monotonically decreasing, i.e., the slope $1/C_{s}$ in fig. 1 monotonically increases, eq. 10 would be violated because concave $\log n_{B}(\log\mu_{B})$ functions cannot surpass the mean value EOS. By contradiction, eq. 10 necessitates a minimum preceding a maximum.
@@
when $C_{s,\mathrm{mean}}^{\mathrm{UD}}<1/3$. A proof is presented in appendix A. Intuitively, eq. 12 assures a second peak because EOSs obeying it are soft in the ultra-dense regime, resembling the maximally soft model characterized by a low trough preceding a high peak.
@@ III Bayesian evidence
Table 1 shows Bayes evidence for scenarios where two peaks are required. It is sensitive to the pQCD renormalization scale $X=\bar{\Lambda}/(\mu_{B}/3)$ and reveals diametric predilections in the extreme cases $X=1,4$. Larger $X$’s lend stronger support because they predict higher $P_{\text{pQCD}}$ thus higher $\Delta P$. Marginalization over this uncertainty is achieved by sampling X from a uniform distribution over the commonly assumed range $X\in[1,4]$ and more restrictive priors motivated by the fact that $P_{\text{pQCD}}$ near $X=1$ is susceptible to pairing contributions [64], by resummed leading- and next-to-leading-order soft contributions [55, 59], and by hints from lattice calculations at finite isospin [33, 34, 86]. The results generally favor the two-peak scenario, and are predominantly determined by the NS maximum mass $M_{\text{TOV}}$ where Bayes factors grow rapidly with increasing $M_{\text{TOV}}$. Constraints on NS radii have very limited effects because current uncertainties are large, and that NS radii may not robustly inform the EOS near maximum mass except in extreme scenarios [64]. Gravitational wave constraints on tidal deformability $\Lambda_{1.4M_{\odot}}\lesssim 600-800$ are implied by the N3LO ${\chi\text{EFT}}$ EOS up to $2n_{0}$ [49, 64]. Results employing ${\chi\text{EFT}}$ to lower densities are reported in appendix C.
@@
Figure 2 shows constraints on the magnitude of the peak and the trough. While posteriors on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ only marginally deviate from the prior suggesting limited constraints on this upper bound, the minimum sound speed is informed. For $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, one finds ${\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}}=0.05^{+0.08}_{-0.04}$ at the $90\%$ level. Considering the prior is $0.16^{+0.15}_{-0.13}$, this presents a strong case for first-order phase transitions above NS densities. This evidence becomes stronger when the second peak is required (shaded). Furthermore, it is interesting to note that among the subset of samples disfavoring low $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$, the preference for the two-peak scenario grows because $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases with increasing $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$. The strengthened evidence in the absence of phase transitions is listed in appendix C. Finally, the onset density [66] of the second peak is inferred to lie in the range $n_{B}\in[4.5_{-0.3}^{+0.5},24^{+3}_{-6}]n_{0}$ or $\mu_{B}\in[1.8^{+0.1}_{-0.2},1.9^{+0.2}_{-0.1}]$ GeV when $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, and $n_{B}\in[5.8_{-0.9}^{+1.0},26^{+4}_{-9}]n_{0}$, or $\mu_{B}\in[1.7^{+0.2}_{-0.2},2.0^{+0.3}_{-0.3}]$ GeV if $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. Detailed examination of the magnitude and location of extrema of $C_{s}$ will be reported in a subsequent paper.
@@ IV Discussion
The second peak above NS densities arises because of asymptotic freedom of QCD and indications that EOS in the NS core is stiff. Since the weakly-coupled CQM is almost conformal  [^3] One may not expect large sensitivities towards $X$ considering $C_{s}^{\text{pQCD}}\simeq 1/3$ regardless of its value. But the pressure $P(\mu_{B})$ is an integrated quantity (eqs. 8 and 4) and tiny deviations in $C_{s}$ originated at asymptotic densities cumulates to a considerable spread down to $\mu_{B}\lesssim 3$ GeV. , a featureless and mostly constant $C_{s}=C_{s,\mathrm{mean}}^{\mathrm{UD}}\simeq 1/3$ above NS densities is only feasible if the EOS is also conformal at the center of most massive NSs. Astrophysical observations suggest rapid stiffening of the EOS above $\sim 2-5n_{0}$ (e.g., [39, 7]), yielding relatively low pressure at $\mu_{B}\simeq 1.5-1.9$ GeV. Taken together with a high $P_{\text{pQCD}}$, a minimum $C_{s,\mathrm{min}}^{\mathrm{UD}}<1/3$ preceding a maximum $C_{s,\mathrm{max}}^{\mathrm{UD}}>1/3$ is thus required to bridge a large $\Delta P=P_{\mathrm{CQM}}-P_{\text{TOV}}$. A related measure is the trace anomaly $P/\mathcal{E}-1/3$ whose density evolution informs structures in $C_{s}$  [90]. Note though successful matching of the $P(\mathcal{E})$ relation are not always consistent with thermodynamics as $\mu_{B}$ could be discontinuous.
@@
Non-perturbative effects in CQM are strongly indicated by the second peak near the weak coupling regime because perturbative calculations generally reveal $C_{s}^{\mathrm{CQM}}<1/3$ [50, 51, 52, 53, 54, 55, 56, 57, 58, 55]. Color superconductivity is perhaps the most anticipated mechanism. Since pairing contributions to the sound speed
$C_{s}^{\Delta}\sim(\Delta(\mu_{B})/\mu_{B})^{2}+\Delta^{\prime}(\mu_{B})\Delta(\mu_{B})/\mu_{B}$
@@
where $\Delta(\mu_{B})$ is the superconducting gap, a maximum exceeding $1/3$ on top of $C_{s}^{\text{pQCD}}\simeq 0.31-0.32$ suggests sizable gaps with non-trivial density dependencies. While estimates based on both leading-order pQCD and phenomenological models are consistent with a decreasing $C^{\Delta}_{s}$ with increasing $\mu_{B}$, only the latter predicts sufficiently large gaps $\Delta\gtrsim 100~{}\text{MeV}$ at $\mu_{B}\simeq 1.6-2.4~{}\text{GeV}$. Therefore, if superconducting gap is the underlying mechanism, it likely receives sizable instanton-induced enhancements [91, 18] or is underestimated by leading-order pQCD calculations in the not-so-weakly-coupled regime. Alternatively, the second peak may be attributed to transitions between asymmetric superconducting phases which give rise to a sharp $\Delta^{\prime}(\mu_{B})$ (though the $\Delta/\mu_{B}$ suppression remains), or meson condensates [92, 93]. These possibilities will be investigated in detail later.
@@
Although Bayesian analyses (e.g., [12, 90] and the present work) generally support violations of the conformal limit [94, 95, 46, 96, 97, 98, 99, 100, 101], $C_{s,\mathrm{max}}^{\mathrm{NS}}\leq 1/3$ is yet to be ruled out by current astrophysical data. This possibility is nevertheless tightly constrained [48, 102, 103], or from a Bayesian perspective highly fine-tuned hence penalized by Occam’s factors implicit in Bayesian inferences. A fully model-independent assessment of the two-peak evidence eschewing parametrization of the NS EOS and statistical interpretations will be reported later.
@@ Acknowledgment
Among the two maxima (minima), at least one of them is the global maxima (minima) across the entire ultra-dense phase $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ ($C_{s,\mathrm{min}}^{\mathrm{UD*}}$).
Next, I show that $C_{s,\mathrm{mean}}^{\text{I}}$ and $C_{s,\mathrm{mean}}^{\text{II}}$ are themselves constrained. These bounds can be obtained by comparing the slopes of these EOSs in fig. 3 to known values and recalling that the inverse of the slopes are the speed of sound squared (see eq. 4). Since the mean value EOS in region I is steeper than the mean value EOS over the entire range (underlying $C_{s,\mathrm{mean}}^{\mathrm{UD}}$), and the mean value EOS in region II is shallower than the red segment with constant $C_{s,\mathrm{max}}^{\mathrm{UD}}=1/3$, one finds
@@
Figure 4: Proving eq. 10 guarantees two peaks when $C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$. The region indicated by eq. 10 is shaded in red, from which I arbitrarily picked a point labeled “A”. Note that in fig. 1 this region is shown as the blue hatches. Here, the maximally soft EOS with $C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$ is infeasible because $C_{s,\mathrm{max}}^{\mathrm{UD*}}>C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$.
@@
Upon substituting in eq. 15 one recovers eq. 19. Even though this scenario with $C_{s,\mathrm{min}}^{\mathrm{UD}}>1/3$ is disfavored by the existence of two-solar-mass NSs, it is still included because (i) in the Bayesian framework eventually excluded samples still require a physical likelihood to be evaluated; and (ii) it is relevant if ${\chi\text{EFT}}$ breaks down earlier, for instance due to a low-density quarkyonic transition. Analyses that follow ${\chi\text{EFT}}$ to lower densities are presented in appendix C.
Appendix B Agnostic EOS models
This appendix presents a family of sound speed based parameterizations used to describe the poorly-understood inner cores of cold NSs. Parameterizing the sound speed, the slope of the EOS, instead of the EOS itself is convenient because physical constraints can be easily imposed. $C_{s}$ is bounded from below by stability (equivalent to the convexity requirement) as $C_{s}>0$ and above by causality as $C_{s}<1$. I initially developed this ensemble in [104] and later incorporated developments in ref  [85]. Some subsets have also been reported elsewhere in e.g. [47, 105, 106]. Note that these parameterizations are only used to describe the beta-equilibrium matter in neutron star interiors where each EOS is truncated at the central density of the heaviest neutron stars i.e., the TOV point.
@@ Appendix B Agnostic EOS models
In principal, $C_{s}$ may be parameterized in terms of any thermodynamic variables at zero temperature, i.e. pressure $P$, energy density $\mathcal{E}$, baryon number density $n$, and baryon chemical potential $\mu$. Zero-temperature thermodynamics ensures that knowledge of the relation between any pair would uniquely determine the rest (except when $C_{s}=0$, see discussions below). These possibilities and the corresponding differential equations are
@@
As a technical side note, when splines are used in conjunction with the option $C_{s}(\mathcal{E})$, closed form integrations exist, providing a test for numerical accuracy. I adopt an embedded 8th order Rugge-Kutta solver and have verified that my implementation with a local error tolerance $\epsilon\sim 10^{-9}$ performs better than $10^{-5}$ globally.
@@ B.3 Interpolation Methods
    The hyper-parameters $l_{\mathrm{min}}$ and $l_{\mathrm{max}}$ are themselves randomly selected from $\mathcal{O}(10^{5})$ priors each, either uniformly or from log-uniform distributions. The resulting normalized length scales $\tilde{l}\equiv l/x_{\mathrm{max}}$ covers a wide range from $1/500$ to $2$. Each choice of $l_{\mathrm{min}}$ and $l_{\mathrm{max}}$ are used to generate about a thousand EOSs, yielding $\mathcal{O}(10^{8})$ samples in total.
@@
    An implicit assumption in GP is density-independent correlation lengths and strengths implied by the translational-invariant kernels $K(x,x^{\prime})=K(|x-x^{\prime}|)$. Ref  [85] used GP to model ${\chi\text{EFT}}$ predictions of the energy per particle where the authors carefully examined the inter-density correlations to ensure a translational invariant kernel over the Fermi momentum is faithful. Density-independent correlations in $C_{s}$ throughout neutron star densities are yet to be justified. This issue is especially severe when $C_{s}$ is low. As mentioned earlier in section B.1, first-order phase transitions introduce tiny length scales in $P$ and $\mu$ whereas in $\mathcal{E}$ and $n$ they manifest as large jumps. As such, translational invariant correlations limit the possible behaviors of generated EOSs  [^4] To be clear, in the continuum and infinite sample size limits, reasonable implementations of GP models and other parameterizations are expected to cover the entire physical phase space. The rate of convergence and the shape of distributions however are sensitive to these choices [102].. A simple remedy is to include all possible $C_{s}(x)$ discussed in section B.1, so that the mixture contains a wider variety of behaviors of $C_{s}$.
@@
Choices of priors are perhaps the most significant factor influencing posteriors in Bayesian analyses. The abscissae of the knots are controlled by the length scale $l\equiv\Delta x$, and its prior is discussed in the previous section. Priors on $C_{s}$ can either be imposed directly or indirectly through priors on $y$. Figure 5 shows the probability density function (pdf) of $C_{s}$ assuming a standard normal distribution on $y\sim\mathcal{N}(0,1)$. Note that the standard deviation of the prior on $y$ is degenerate with the scaling of $y$ in the mappings, which is set to $1$ in table 2. The choice of mappings implicitly determines the shape of pdf of $C_{s}$. I have included all four options in the ensemble, and have also repeated the above steps but impose priors on $c_{s}=\sqrt{C_{s}}$ instead of $C_{s}$. Compared to $C_{s}$-based models, directly working with $\sqrt{C_{s}}$ generally favors lower sound speed in the prior, and sometimes notable differences in the posteriors are observed between the choices of $C_{s}$ and $\sqrt{C_{s}}$ [102].

### v3

@@ I Introduction
It can be understood in terms of the geometric interpretation of eq. 8, namely $\Delta P$ as the area in fig. 1. Continuations of NS EOSs satisfying eq. 10 must have segments above the mean value EOS. If the sound speed of such high-density models were to decrease monotonically, i.e., the slope $1/C_{s}$ in fig. 1 were to monotonically increase, eq. 10 would be violated because convex $\log n_{B}(\log\mu_{B})$ functions cannot surpass the mean value EOS in fig. 1. By contradiction, eq. 10 necessitates a minimum preceding a maximum.
@@
when $C_{s,\mathrm{mean}}^{\mathrm{UD}}<1/3$. A proof is presented in appendix A. Intuitively, eq. 12 assures a second peak because EOSs obeying it are soft in the ultra-dense regime and must have segments resembling the maximally soft model characterized by a low trough preceding a high peak.
@@ III Bayesian evidence
Table 1 shows Bayes evidence for scenarios where two peaks are required. It is sensitive to the pQCD renormalization scale $X=\bar{\Lambda}/(\mu_{B}/3)$ and reveals diametric predilections in the extreme cases $X=1,4$. Larger $X$’s lend stronger support because they predict higher $P_{\text{pQCD}}$ thus higher $\Delta P$. Marginalization over this uncertainty is achieved by sampling X from a uniform distribution over the commonly assumed range $X\in[1,4]$ and more restrictive priors motivated by the fact that $P_{\text{pQCD}}$ near $X=1$ is susceptible to pairing contributions [64], by resummed leading and next-to-leading soft contributions [55, 59], and by hints from lattice calculations at finite isospin [86, 33, 34]. The results generally favor the two-peak scenario, and are predominantly determined by the NS maximum mass $M_{\text{TOV}}$ where Bayes factors grow rapidly with increasing $M_{\text{TOV}}$. Constraints on NS radii have very limited effects because current uncertainties are large, and that NS radii may not robustly inform the EOS near maximum mass except in extreme scenarios [64]. Gravitational wave constraints on tidal deformability $\Lambda_{1.4M_{\odot}}\lesssim 600-800$ are implied by the N3LO ${\chi\text{EFT}}$ EOS up to $2n_{0}$ [49, 64]. Results employing ${\chi\text{EFT}}$ to lower densities are reported in appendix C.
@@
Figure 2 shows constraints on the magnitude of the peak and the trough. While posteriors on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ only marginally deviate from the prior suggesting limited constraints on this upper bound, the minimum sound speed is informed. For $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, one finds ${\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}}=0.05^{+0.08}_{-0.04}$ at the $90\%$ level. Considering the prior is $0.16^{+0.15}_{-0.13}$, this presents a strong case for first-order phase transitions above NS densities. This evidence becomes stronger when the second peak is required (shaded). Furthermore, it is interesting to note that among the subset of samples disfavoring low $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$, the preference for the two-peak scenario grows because $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases with increasing $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$. The strengthened evidence in the absence of phase transitions is listed in appendix C. Finally, the onset density [66] of the second peak is inferred to lie in the range $n_{B}\in[4.5_{-0.3}^{+0.5},24^{+3}_{-6}]n_{0}$ or $\mu_{B}\in[1.8^{+0.1}_{-0.2},1.9^{+0.2}_{-0.1}]$ GeV when $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, and $n_{B}\in[5.8_{-0.9}^{+1.0},26^{+4}_{-9}]n_{0}$ or $\mu_{B}\in[1.7^{+0.2}_{-0.2},2.0^{+0.3}_{-0.3}]$ GeV if $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. Detailed examination of the magnitude and location of extrema of $C_{s}$ will be reported in a subsequent paper.
@@ IV Discussion
The second peak above NS densities arises because of asymptotic freedom of QCD and indications that EOS in the NS core is stiff. Since the weakly-coupled CQM is almost conformal  [^3] One may not expect large sensitivities towards $X$ considering $C_{s}^{\text{pQCD}}\simeq 1/3$ regardless of its value. But the pressure $P(\mu_{B})$ is an integrated quantity (eqs. 8 and 4) and tiny deviations in $C_{s}$ originated at asymptotic densities cumulate to a considerable spread down to $\mu_{B}\lesssim 3$ GeV. , a featureless and mostly constant $C_{s}=C_{s,\mathrm{mean}}^{\mathrm{UD}}\simeq 1/3$ above NS densities is only feasible if the EOS is also conformal at the center of most massive NSs. Astrophysical observations suggest rapid stiffening of the EOS above $\sim 2-5n_{0}$ (e.g., [39, 7]), yielding relatively low pressure at $\mu_{B}\simeq 1.5-1.9$ GeV. Taken together with a high $P_{\text{pQCD}}$, a minimum $C_{s,\mathrm{min}}^{\mathrm{UD}}<1/3$ preceding a maximum $C_{s,\mathrm{max}}^{\mathrm{UD}}>1/3$ is thus required to bridge a large $\Delta P=P_{\mathrm{CQM}}-P_{\text{TOV}}$. These implications of stiff NS cores are captured by eqs. 10 and 12 rooted in the continuity of $P(\mu_{B})$. A related measure is the trace anomaly $P/\mathcal{E}-1/3$ whose density evolution informs structures in $C_{s}$  [90]. Note though successful matching of the $P(\mathcal{E})$ relation is not always consistent with thermodynamics as $\mu_{B}$ could be discontinuous.
@@
Non-perturbative effects in CQM are strongly indicated by the second peak near the weak-coupling regime because perturbative calculations generally reveal $C_{s}^{\mathrm{CQM}}<1/3$ [50, 51, 52, 53, 54, 55, 56, 57, 58, 55]. Color superconductivity is perhaps the most anticipated mechanism. Since pairing contributions to the sound speed
$C_{s}^{\Delta}\approx 4\frac{\Delta^{2}}{\mu_{B}^{2}}-2\left[{\Delta^{\prime}}^{2}+\frac{\Delta\Delta^{\prime}}{\mu_{B}}\right]+\mathcal{O}(\Delta^{\prime 2},(\frac{\Delta}{\mu_{B}})^{2},\Delta^{(n>1)})$
@@
where $\Delta=\Delta(\mu_{B})$ is the superconducting gap, a maximum exceeding $1/3$ on top of $C_{s}^{\text{pQCD}}\simeq 0.31-0.32$ suggests sizable gaps. While estimates based on both leading-order pQCD and phenomenological models yield the postulated density dependence for $C_{s}^{\mathrm{CQM}}=C_{s}^{\text{pQCD}}+C_{s}^{\Delta}$ since $\Delta\sim\exp(-1/g_{s})$ falls off faster than any powers of $g_{s}$ in $C_{s}^{\text{pQCD}}=1/3+\mathcal{O}(g_{s}^{2})$ as $\mu_{B}\rightarrow\infty$, only the latter predicts sufficiently large gaps $\Delta\gtrsim 100~{}\text{MeV}$ at $\mu_{B}\simeq 1.6-2.4~{}\text{GeV}$. Therefore, if superconducting gap is the underlying mechanism, it likely receives sizable instanton enhancements [91, 18] or is underestimated by leading-order pQCD calculations in the not-so-weakly-coupled regime. Alternatively, the second peak may be attributed to meson condensates [92, 93]. These possibilities will be investigated in detail later.
@@
Although Bayesian analyses (e.g., [12, 90] and the present work) generally support violations of the conformal limit [94, 95, 46, 96, 97, 98, 99, 100, 101] in NSs, $C_{s,\mathrm{max}}^{\mathrm{NS}}\leq 1/3$ is yet to be ruled out by current astrophysical data. This possibility is nevertheless tightly constrained [48, 102, 103], or from a Bayesian perspective highly fine-tuned hence penalized by Occam’s factors implicit in Bayesian inferences. A fully model-independent assessment of the two-peak evidence eschewing parametrization of the NS EOS and statistical interpretations will be reported later.
@@ Acknowledgment
Among the two maxima (minima), at least one of them is the global maximum (minimum) across the entire ultra-dense phase $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ ($C_{s,\mathrm{min}}^{\mathrm{UD*}}$).
Next, I show that $C_{s,\mathrm{mean}}^{\text{I}}$ and $C_{s,\mathrm{mean}}^{\text{II}}$ are themselves constrained. These bounds can be obtained by comparing the slopes of these EOSs in fig. 3 to known values and recalling that the inverse of the slopes are the speed of sound squared (see eq. 4). Since the mean value EOS in region I is steeper than the mean value EOS over the entire range (underlying $C_{s,\mathrm{mean}}^{\mathrm{UD}}$), and the mean value EOS in region II is shallower than the red segment with constant $C_{s}=C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$, one finds
@@
Figure 4: Proving eq. 10 guarantees two peaks when $C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$. The region indicated by eq. 10 is shaded in red, from which I arbitrarily picked a point labeled “A”. Note that in fig. 1 this region is shown as the blue hatches. Here, the maximally soft EOS with $C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3$ is infeasible because $C_{s,\mathrm{max}}^{\mathrm{UD*}}\geq C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$.
@@
Upon substituting in eq. 15 one recovers eq. 19. Even though this scenario where $C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$ is disfavored by the existence of two-solar-mass NSs, it is still included because (i) in the Bayesian framework eventually excluded samples still require a physical likelihood to be evaluated; and (ii) it is relevant if ${\chi\text{EFT}}$ breaks down earlier, for instance due to a low-density quarkyonic transition. Analyses that follow ${\chi\text{EFT}}$ to lower densities are presented in appendix C.
Appendix B Agnostic models of NS cores
This appendix presents a family of sound speed based parameterizations used to describe the poorly-understood inner cores of cold NSs. Parameterizing the sound speed, the slope of the EOS, instead of the EOS itself is convenient because physical constraints can be easily imposed. $C_{s}$ is bounded from below by stability (equivalent to the convexity requirement) as $C_{s}>0$ and above by causality as $C_{s}<1$. I initially developed this ensemble in [104] and later incorporated developments in ref  [105, 85]. Some subsets have also been reported elsewhere in e.g. [47, 106]. Note that these parameterizations are only used to describe the beta-equilibrium matter in neutron star interiors where each EOS is truncated at the central density of the heaviest neutron stars i.e., the TOV point.
@@ Appendix B Agnostic models of NS cores
In principal, $C_{s}$ may be parameterized in terms of any thermodynamic variables at zero temperature, i.e. pressure $P$, energy density $\mathcal{E}$, baryon number density $n$, and baryon chemical potential $\mu=(P+\mathcal{E})/n$. Zero-temperature thermodynamics ensures that knowledge of the relation between any pair would uniquely determine the rest (except when $C_{s}=0$, see discussions below). These possibilities and the corresponding differential equations are
@@
As a technical side note, when splines are used in conjunction with the option $C_{s}(\mathcal{E})$, closed form integrations exist, providing a test for numerical accuracy. I adopt an embedded 8th-order Rugge-Kutta routine [107] and have verified that my implementation with a local error tolerance $\epsilon\sim 10^{-9}$ performs better than $10^{-5}$ globally.
@@ B.3 Interpolation Methods
    The hyper-parameters $l_{\mathrm{min}}$ and $l_{\mathrm{max}}$ are themselves randomly selected from $\mathcal{O}(10^{5})$ priors each, either uniformly or from log-uniform distributions. The resulting normalized length scales $\tilde{l}\equiv l/(x_{\mathrm{max}}-x_{\chi\text{EFT}})$ covers a wide range from $1/2000$ to $2$, where $x_{\chi\text{EFT}}$ refers to the value of $x$ at $n_{B}=n_{\chi\text{EFT}}$. Each choice of $l_{\mathrm{min}}$ and $l_{\mathrm{max}}$ are used to generate about a thousand EOSs, yielding $\mathcal{O}(10^{8})$ samples in total.
@@
    An implicit assumption in GP is density-independent correlation lengths and strengths implied by the translational-invariant kernels $K(x,x^{\prime})=K(|x-x^{\prime}|)$. Ref  [85] used GP to model ${\chi\text{EFT}}$ predictions of the energy per particle where the authors carefully examined the inter-density correlations to ensure a translational invariant kernel over the Fermi momentum is faithful. Density-independent correlations in $C_{s}$ throughout neutron star densities are yet to be justified. This issue is especially notable when $C_{s}$ is low. As mentioned earlier in section B.1, first-order phase transitions introduce tiny length scales in $P$ and $\mu$ whereas in $\mathcal{E}$ and $n$ they manifest as large jumps. As such, translational invariant correlations limit the possible behaviors of generated EOSs  [^4] To be clear, in the continuum and infinite sample size limits, reasonable implementations of GP models and other parameterizations are expected to cover the entire physical phase space. The rate of convergence and the shape of probability distributions however are sensitive to these choices [102].. A simple remedy is to include all possible $C_{s}(x)$ discussed in section B.1, so that the mixture contains a wider variety of behaviors of $C_{s}$.
@@
Choices of priors are perhaps the most significant factor influencing posteriors in Bayesian analyses. The abscissae of the knots are controlled by the length scale $l\equiv\Delta x$, and its prior is discussed in the previous section. Priors on $C_{s}$ can either be imposed directly or indirectly through priors on $y$. Figure 5 shows the probability density function (pdf) of $C_{s}$ assuming a standard normal distribution on $y\sim\mathcal{N}(0,1)$. Note that the standard deviation of the prior on $y$ is degenerate with the scaling of $y$ in the mappings, which is set to $1$ in table 2. The choice of mappings implicitly determines the shape of pdf of $C_{s}$. I have included all four options in the ensemble, and have also repeated the above steps but impose priors on $c_{s}=\sqrt{C_{s}}$ instead of $C_{s}$. Compared to $C_{s}$-based models, directly working with $\sqrt{C_{s}}$ favors lower sound speed in the prior, and sometimes notable differences in the posteriors are observed between the choices of $C_{s}$ and $\sqrt{C_{s}}$ [102].